# Optimal rates of approximation by shallow ReLUk neural networks and applications to nonparametric regression

### 21 Jun 2023 v2

@@ 1 Introduction
Neural networks generate very popular function classes used in machine learning algorithms [Anthony and Bartlett, 2009; LeCun et al., 2015]. The fundamental building blocks of neural networks are ridge functions (also called neurons) of the form $x\in\mathbb{R}^{d}\mapsto\rho((x^{\intercal},1)v)$, where $\rho:\mathbb{R}\to\mathbb{R}$ is a continuous activation function and $v\in\mathbb{R}^{d+1}$ is a trainable parameter. It is well-known that a shallow neural network with non-polynomial activation
@@
is universal in the sense that it can approximate any continuous function on any compact set with desired accuracy when the number of neurons $N$ is sufficiently large [Cybenko, 1989; Hornik, 1991; Pinkus, 1999]. The approximation and statistical properties of neural networks with different architectures have also been widely studied in the literature [Yarotsky, 2017; Zhou, 2020b; Bartlett et al., 2019; Schmidt-Hieber, 2020], especially when $\rho$ is a sigmoidal activation or $\rho$ is the ReLUk function $\sigma_{k}(t)=\max\{0,t\}^{k}$, the $k$-power of the rectified linear unit (ReLU) with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
The main focus of this paper is rates of approximation by neural networks. For classical smooth function classes, such as Hölder functions, Mhaskar [1996] (see also [Pinkus, 1999, Theorem 6.8]) presented the optimal approximation rates for shallow neural networks, when the activation function $\rho\in C^{\infty}(\Omega)$ is not a polynomial on some open interval $\Omega$ (ReLUk does not satisfy this condition). Similar results have been established for deep ReLU neural networks in recent works [Yarotsky, 2017, 2018; Shen et al., 2020; Lu et al., 2021]. All these approximation rates are obtained by using the idea that one can construct neural networks to approximate polynomials efficiently. There is another line of works [Barron, 1993; Makovoz, 1996; Klusowski and Barron, 2018; Siegel and Xu, 2020, 2022] studying the approximation rates for functions of certain integral forms (such as (1.2)) by using a random sampling argument due to Maurey [Pisier, 1981]. In particular, Barron [1993] derived dimension independent approximation rate for sigmoid type activations and functions $h$, whose Fourier transform $\widehat{h}$ satisfies $\int_{\mathbb{R}^{d}}|\omega||\widehat{h}(\omega)|d\omega<\infty$. This result has been improved and generalized to ReLU in recent articles [Klusowski and Barron, 2018; Siegel and Xu, 2020, 2022].
In this paper, we give a connection of these two lines of approximation results for neural networks (i.e. the constructive approximation of smooth functions and the random approximation of integral representation). Specifically, we consider the following function class defined on the unit ball $\mathbb{B}^{d}$ of $\mathbb{R}^{d}$ induced by vectors on unit sphere $\mathbb{S}^{d}$ of $\mathbb{R}^{d+1}$ as
@@
which can be regarded as a shallow ReLUk neural network with infinite width [Bach, 2017]. The variation norm restriction $\|\mu\|\leq M$ gives a constraint on the size of the weights in the network. We study how well $\mathcal{F}_{\sigma_{k}}(M)$ approximates the unit ball of Hölder class $\mathcal{H}^{\alpha}$ with smoothness index $\alpha>0$ as $M\to\infty$. Roughly speaking, our main theorem shows that, if $\alpha>(d+2k+1)/2$, then $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ depending on $k,d,\alpha$, and if $\alpha<(d+2k+1)/2$, we obtain the approximation bound
@@
where, for two quantities $X$ and $Y$, $X\lesssim Y$ (or $Y\gtrsim X$) denotes the statement that $X\leq CY$ for some constant $C>0$ (we will also denote $X\asymp Y$ when $X\lesssim Y\lesssim X$). In other words, sufficiently smooth functions are always contained in the shallow neural network space $\mathcal{F}_{\sigma_{k}}(M)$. And, for less smooth functions, we can characterize the approximation error by the variation norm. Furthermore, combining our result with the random approximation bounds from Bach [2017]; Siegel and Xu [2022], we are able to prove that shallow ReLUk neural network of the form (1.1) achieves the optimal approximation rate $\mathcal{O}(N^{-\alpha/d})$ for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, which generalizes the result of [Mhaskar, 1996] to ReLUk.
In addition to shallow neural networks, we can also apply our results to derive approximation bounds for multi-layer neural networks and convolutional neural networks (CNNs) when $k=1$ (ReLU activation $\sigma:=\sigma_{1}$). These approximation bounds can then be used to study the performances of machine learning algorithms based on neural networks [Anthony and Bartlett, 2009]. Here, we illustrate the idea by studying the nonparametric regression problem. The goal is to learn a function $h$ from a hypothesis space $\mathcal{H}$ from its noisy samples
@@
where $\mathcal{F}_{n}$ is an appropriately chosen function class. For instance, in deep learning, $\mathcal{F}_{n}$ is parameterized by deep neural networks and one solves the minimization by (stochastic) gradient descent methods. Assuming that we can compute a minimizer $f_{n}^{*}\in\mathcal{F}_{n}$, the performance of the algorithm is often measured by the square loss $\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}$. A fundamental question in learning theory is to determine the convergence rate of the error $\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\to 0$ as the sample size $n\to\infty$. The error can be decomposed into two components: approximation error and generalization error (also called estimation error). For neural network models $\mathcal{F}_{n}$, our results provide bounds for the approximation errors, while the generalization errors can be bounded by the complexity of the models [Shalev-Shwartz and Ben-David, 2014; Mohri et al., 2018]. We study the cases $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$ for three ReLU neural network models: shallow neural network, over-parameterized neural network, and CNN. The models and our contributions are summarized as follows:
@@
    Shallow ReLU neural network $\mathcal{F}_{\sigma}(N,M)$, where $N$ is the number of neurons and $M$ is a bound for the variation norm that measures the size of the weights. We prove optimal approximation rates (in terms of $N$) for this model. It is also shown that this model can achieve the optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$, which complements the recent results for deep neural networks [Schmidt-Hieber, 2020; Kohler and Langer, 2021].
@@
    Over-parameterized (deep or shallow) ReLU neural network $\mathcal{NN}(W,L,M)$ studied in [Jiao et al., 2023], where $W,L$ are the width and depth respectively, and $M$ is a constraint on the weight matrices. For fixed depth $L$, the generalization error for this model can be controlled by $M$ [Jiao et al., 2023]. When $\mathcal{H}=\mathcal{H}^{\alpha}$, we characterize the approximation error by $M$, and allow the width $W$ to be arbitrary large so that the model can be over-parameterized (the number of parameters is larger than the number of samples). When $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we can simply increase the width to reduce the approximation error so that the model can also be over-parameterized. Our result shows that this model can achieve nearly optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$. Both the approximation and convergence rates improve the results of [Jiao et al., 2023].
@@
    Sparse convolutional ReLU neural network $\mathcal{CNN}(s,L)$ introduced by Zhou [2020b], where $L$ is the depth and $s\geq 2$ is a fixed integer that controls the filter length. This model is shown to be universal for approximation [Zhou, 2020b] and universal consistent for regression [Lin et al., 2022]. We improve the approximation bound in [Zhou, 2020b] and give new convergence rates of this model for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$.
@@
The rest of the paper is organized as follows. In Section 2, we present our approximation results for shallow neural networks. Section 3 gives a proof of our main theorem. In Section 4, we apply our approximation results to study there neural network models and derive convergence rates for nonparametric regression using these models. Section 5 concludes this paper with a discussion on possible future directions of research.
@@ 2 Approximation rates for shallow neural networks
where $\mu$ is a signed Radon measure on $\mathbb{S}^{d}$ with finite total variation $\|\mu\|:=|\mu|(\mathbb{S}^{d})<\infty$ and $\sigma_{k}(t):=\max\{t,0\}^{k}$ with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ is the ReLUk function (when $k=0$, $\sigma_{0}(t)$ is the Heaviside function). For simplicity, we will also denote the ReLU function by $\sigma:=\sigma_{1}$. The variation norm $\gamma(f)$ of $f$ is the infimum of $\|\mu\|$ over all decompositions of $f$ as (2.1) [Bach, 2017]. By the compactness of $\mathbb{S}^{d}$, the infimum can be attained by some signed measure $\mu$. We denote $\mathcal{F}_{\sigma_{k}}(M)$ as the function class that contains all functions $f$ in the form (2.1) whose variation norm $\gamma(f)\leq M$, see (1.2). The class $\mathcal{F}_{\sigma_{k}}(M)$ can be thought of as an infinitely wide neural network with a constraint on its weights. The variation spaces corresponding to shallow neural networks have been studied by many researchers. We refer the reader to [Savarese et al., 2019; Ongie et al., 2020; Bartolucci et al., 2023; Parhi and Nowak, 2022, 2023; Siegel and Xu, 2020, 2022, 2023] for several other definitions and characterizations of these spaces.
@@
where $s=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is an multi-index and
$\|f\|_{C^{r}(\mathbb{R}^{d})}:=\max_{\|s\|_{1}\leq r}\|\partial^{s}f\|_{L^{\infty}(\mathbb{R}^{d})},\quad|f|_{C^{0,\beta}(\mathbb{R}^{d})}:=\sup_{x\neq y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{\|x-y\|_{2}^{\beta}},$
where we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm, since we mainly consider continuous functions. We write $C^{r,\beta}(\mathbb{B}^{d})$ for the Banach space of all restrictions to $\mathbb{B}^{d}$ of functions in $C^{r,\beta}(\mathbb{R}^{d})$. The norm is given by $\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}=\inf\{\|g\|_{C^{r,\beta}(\mathbb{R}^{d})}:g\in C^{r,\beta}(\mathbb{R}^{d})\mbox{ and }g=f\mbox{ on }\mathbb{B}^{d}\}$. For convenience, we will denote the unit ball of $C^{r,\beta}(\mathbb{B}^{d})$ by $\mathcal{H}^{\alpha}:=\{f\in C^{r,\beta}(\mathbb{B}^{d}):\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}\leq 1\}$. Note that, for $\alpha=1$, $\mathcal{H}^{\alpha}$ is Lipschitz continuous function class.
Due to the universality of shallow neural networks [Pinkus, 1999], $\mathcal{F}_{\sigma_{k}}(M)$ can approximate any continuous functions on $\mathbb{B}^{d}$ if $M$ is sufficiently large. Our main theorem estimates the rate of this approximation for Hölder class.
@@
The proof of Theorem 2.1 is deferred to the next section. Our proof uses similar ideas as Bach [2017, Proposition 3], which obtained the same approximation rate for $\alpha=1$. The conclusion is more complicated for the critical value $\alpha=(d+2k+1)/2$. We think this is due to the proof technique and conjecture that $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for all $\alpha\geq(d+2k+1)/2$, see Remark 3.4. Nevertheless, in practical applications of machine learning, the dimension $d$ is large and it is reasonable to expect that $\alpha<(d+2k+1)/2$.
@@
where we restrict the measure $\mu$ to be a discrete one supported on at most $N$ points. The next proposition shows that any function in $\mathcal{F}_{\sigma_{k}}(M)$ is the limit of functions in $\mathcal{F}_{\sigma_{k}}(N,M)$ as $N\to\infty$.
Proposition 2.2.
Let $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$ denote the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{\infty}(\mathbb{B}^{d})$. If $k\in\mathbb{N}$, then $\mathcal{F}_{\sigma_{k}}(1)=\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$. If $k=0$, then $\mathcal{F}_{\sigma_{0}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{0}}(1)$.
@@
We first show that $\mathcal{F}_{\sigma_{k}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}_{0}$. For any $f\in\mathcal{F}_{\sigma_{k}}(1)$ with the integral form $f(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v)$, we can decompose $f$ as
$f(x)$ $=\|\mu_{+}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{+}\|}-\|\mu_{-}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{-}\|}$
@@
where $\mu_{+}$ and $\mu_{-}$ are the positive and negative parts of $\mu$. If $f_{+},f_{-}\in\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$, then $f\in\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$. Hence, without loss of generality, we can assume $\mu$ is a probability measure. We are going to approximate $f$ by uniform laws of large numbers. Let $\{v_{i}\}_{i=1}^{N}$ be $N$ i.i.d. samples from $\mu$. By symmetrization argument (see Wainwright [2019, Theorem 4.10] for example), we can bound the expected approximation error by Rademacher complexity [Bartlett and Mendelson, 2002]:
@@
where $(\epsilon_{1},\dots,\epsilon_{N})$ is an i.i.d. sequence of Rademacher random variables. For $k\in\mathbb{N}$, the Lipschitz constant of $\sigma_{k}$ on $\sqrt{2},\sqrt{2}]$ is $k2^{(k-1)/2}$. By the contraction property of Rademacher complexity [Ledoux and Talagrand, 1991, Corollary 3.17],
@@
For $k=0$, the VC dimension of the function class $\{f_{x}(v)=\sigma_{0}((x^{\intercal},1)v):x\in\mathbb{B}^{d}\}$ is at most $d$ [Wainwright, 2019, Proposition 4.20]. Thus, we have the bound $\mathcal{E}_{0}(N)\lesssim\sqrt{d/N}$ by Wainwright [2019, Example 5.24]. Hence, $f$ is in the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$.
Next, we show that $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}$. Since $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$, we only need to show that $\mathcal{F}_{\sigma_{k}}(1)$ is closed in $L^{\infty}(\mathbb{B}^{d})$. Let $f_{n}(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu_{n}(v)$, where $\|\mu_{n}\|\leq 1$, be a convergent sequence with limit $f\in L^{\infty}(\mathbb{B}^{d})$. By the compactness of $\mathbb{S}^{d}$ and Prokhorov’s theorem, there exists a weakly convergent subsequence $\mu_{n_{i}}\to\mu$. In particular, $\|\mu\|\leq 1$ and for any $x\in\mathbb{B}^{d}$,

### 09 Jan 2024 v3

@@ 1 Introduction
Neural networks generate very popular function classes used in machine learning algorithms [2, 28]. The fundamental building blocks of neural networks are ridge functions (also called neurons) of the form $x\in\mathbb{R}^{d}\mapsto\rho((x^{\intercal},1)v)$, where $\rho:\mathbb{R}\to\mathbb{R}$ is a continuous activation function and $v\in\mathbb{R}^{d+1}$ is a trainable parameter. It is well-known that a shallow neural network with non-polynomial activation
@@
is universal in the sense that it can approximate any continuous functions on any compact set with desired accuracy when the number of neurons $N$ is sufficiently large [11, 23, 48]. The approximation and statistical properties of neural networks with different architectures have also been widely studied in the literature [64, 67, 6, 52], especially when $\rho$ is a sigmoidal activation or $\rho$ is the ReLUk function $\sigma_{k}(t)=\max\{0,t\}^{k}$, the $k$-power of the rectified linear unit (ReLU) with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
The main focus of this paper is rates of approximation by neural networks. For classical smooth function classes, such as Hölder functions, Mhaskar [39] (see also [48, Theorem 6.8]) presented approximation rates for shallow neural networks, when the activation function $\rho\in C^{\infty}(\Omega)$ is not a polynomial on some open interval $\Omega$ (ReLUk does not satisfy this condition). It is known that the rates obtained by Mhaskar are optimal if the network weights are required to be continuous functions of the target function. Recently, optimal rates of approximation have also been established for deep ReLU neural networks [64, 65, 54, 31], even without the continuity requirement on the network weights. All these approximation rates are obtained by using the idea that one can construct neural networks to approximate polynomials efficiently. There is another line of works [4, 34, 25, 56, 57] studying the approximation rates for functions of certain integral forms (such as (1.2)) by using a random sampling argument due to Maurey [49]. In particular, Barron [4] derived dimension independent approximation rates for sigmoid type activations and functions $h$, whose Fourier transform $\widehat{h}$ satisfies $\int_{\mathbb{R}^{d}}|\omega||\widehat{h}(\omega)|d\omega<\infty$. This result has been improved and generalized to ReLU activation in recent articles [25, 56, 57].
In this paper, we continue the study of these two lines of approximation theories for neural networks (i.e. the constructive approximation of smooth functions and the random approximation of integral representations). Our main result shows how well integral representations corresponding to ReLUk neural networks can approximate smooth functions. By combining this result with the random approximation theory of integral forms, we are able to establish the optimal rates of approximation for shallow ReLUk neural networks. Specifically, we consider the following function class defined on the unit ball $\mathbb{B}^{d}$ of $\mathbb{R}^{d}$ induced by vectors on unit sphere $\mathbb{S}^{d}$ of $\mathbb{R}^{d+1}$ as
@@
which can be regarded as a shallow ReLUk neural network with infinite width [3]. The restriction on the total variation $\|\mu\|:=|\mu|(\mathbb{S}^{d})\leq M$ gives a constraint on the size of the weights in the network. We study how well $\mathcal{F}_{\sigma_{k}}(M)$ approximates the unit ball of Hölder class $\mathcal{H}^{\alpha}$ with smoothness index $\alpha>0$ as $M\to\infty$. Roughly speaking, our main theorem shows that, if $\alpha>(d+2k+1)/2$, then $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ depending on $k,d,\alpha$, and if $\alpha<(d+2k+1)/2$, we obtain the approximation bound
@@
where, for two quantities $X$ and $Y$, $X\lesssim Y$ (or $Y\gtrsim X$) denotes the statement that $X\leq CY$ for some constant $C>0$ (we will also denote $X\asymp Y$ when $X\lesssim Y\lesssim X$). In other words, sufficiently smooth functions are always contained in the shallow neural network space $\mathcal{F}_{\sigma_{k}}(M)$. And, for less smooth functions, we can characterize the approximation error by the variation norm. Furthermore, combining our result with the random approximation bounds from [3, 57, 55], we are able to prove that shallow ReLUk neural network of the form (1.1) achieves the optimal approximation rate $\mathcal{O}(N^{-\alpha/d})$ for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, which generalizes the result of Mhaskar [39] to ReLUk activation.
In addition to shallow neural networks, we can also apply our results to derive approximation bounds for multi-layer neural networks and convolutional neural networks (CNNs) when $k=1$ (ReLU activation $\sigma:=\sigma_{1}$). These approximation bounds can then be used to study the performances of machine learning algorithms based on neural networks [2]. Here, we illustrate the idea by studying the nonparametric regression problem. The goal of this problem is to learn a function $h$ from a hypothesis space $\mathcal{H}$ from its noisy samples
@@
where $\mathcal{F}_{n}$ is an appropriately chosen function class. For instance, in deep learning, $\mathcal{F}_{n}$ is parameterized by deep neural networks and one solves the minimization by (stochastic) gradient descent methods. Assuming that we can compute a minimizer $f_{n}^{*}\in\mathcal{F}_{n}$, the performance of the algorithm is often measured by the square loss $\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}$. A fundamental question in learning theory is to determine the convergence rate of the error $\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\to 0$ as the sample size $n\to\infty$. The error can be decomposed into two components: approximation error and generalization error (also called estimation error). For neural network models $\mathcal{F}_{n}$, our results provide bounds for the approximation errors, while the generalization errors can be bounded by the complexity of the models [53, 40]. We study the cases $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$ for three ReLU neural network models: shallow neural network, over-parameterized neural network, and CNN. The models and our contributions are summarized as follows:
@@
    Shallow ReLU neural network $\mathcal{F}_{\sigma}(N,M)$, where NN is the number of neurons and $M$ is a bound for the variation norm that measures the size of the weights. We prove optimal approximation rates (in terms of NN) for this model. It is also shown that this model can achieve the optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$, which complements the recent results for deep neural networks [52, 27].
@@
    Over-parameterized (deep or shallow) ReLU neural network $\mathcal{NN}(W,L,M)$ studied in [24], where $W,L$ are the width and depth respectively, and $M$ is a constraint on the weight matrices. For fixed depth $L$, the generalization error for this model can be controlled by $M$ [24]. When $\mathcal{H}=\mathcal{H}^{\alpha}$, we characterize the approximation error by $M$, and allow the width $W$ to be arbitrary large so that the model can be over-parameterized (the number of parameters is larger than the number of samples). When $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we can simply increase the width to reduce the approximation error so that the model can also be over-parameterized. Our result shows that this model can achieve nearly optimal convergence rates for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$. Both the approximation and convergence rates improve the results of [24].
@@
    Sparse convolutional ReLU neural network $\mathcal{CNN}(s,L)$ introduced by [67], where $L$ is the depth and $s\geq 2$ is a fixed integer that controls the filter length. This model is shown to be universal for approximation [67] and universal consistent for regression [30]. We improve the approximation bound in [67] and give new convergence rates of this model for learning $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$.
@@
The rest of the paper is organized as follows. In Section 2, we present our approximation results for shallow neural networks. Section 3 gives a proof of our main theorem. In Section 4, we apply our approximation results to study these neural network models and derive convergence rates for nonparametric regression using these models. Section 5 concludes this paper with a discussion on possible future directions of research.
@@ 2 Approximation rates for shallow neural networks
where $\mu$ is a signed Radon measure on $\mathbb{S}^{d}$ with finite total variation $\|\mu\|:=|\mu|(\mathbb{S}^{d})<\infty$ and $\sigma_{k}(t):=\max\{t,0\}^{k}$ with $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ is the ReLUk function (when $k=0$, $\sigma_{0}(t)$ is the Heaviside function). For simplicity, we will also denote the ReLU function by $\sigma:=\sigma_{1}$. The variation norm $\gamma(f)$ of $f$ is the infimum of $\|\mu\|$ over all decompositions of $f$ as (2.1) [3]. By the compactness of $\mathbb{S}^{d}$, the infimum can be attained by some signed measure $\mu$. We denote $\mathcal{F}_{\sigma_{k}}(M)$ as the function class that contains all functions $f$ in the form (2.1) whose variation norm $\gamma(f)\leq M$, see (1.2). The class $\mathcal{F}_{\sigma_{k}}(M)$ can be thought of as an infinitely wide neural network with a constraint on its weights. The variation spaces corresponding to shallow neural networks have been studied by many researchers. We refer the reader to [51, 44, 8, 45, 46, 55, 56, 57, 58] for several other definitions and characterizations of these spaces.
@@
where $s=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index and
$\|f\|_{C^{r}(\mathbb{R}^{d})}:=\max_{\|s\|_{1}\leq r}\|\partial^{s}f\|_{L^{\infty}(\mathbb{R}^{d})},\quad|f|_{C^{0,\beta}(\mathbb{R}^{d})}:=\sup_{x\neq y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{\|x-y\|_{2}^{\beta}}.$
Here we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm, since we mainly consider continuous functions. We write $C^{r,\beta}(\mathbb{B}^{d})$ for the Banach space of all restrictions to $\mathbb{B}^{d}$ of functions in $C^{r,\beta}(\mathbb{R}^{d})$. The norm is given by $\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}=\inf\{\|g\|_{C^{r,\beta}(\mathbb{R}^{d})}:g\in C^{r,\beta}(\mathbb{R}^{d})\mbox{ and }g=f\mbox{ on }\mathbb{B}^{d}\}$. For convenience, we will denote the unit ball of $C^{r,\beta}(\mathbb{B}^{d})$ by
$\mathcal{H}^{\alpha}:=\left\{f\in C^{r,\beta}(\mathbb{B}^{d}):\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}\leq 1\right\}.$
Note that, for $\alpha=1$, $\mathcal{H}^{\alpha}$ is a class of Lipschitz continuous functions.
Due to the universality of shallow neural networks [48], $\mathcal{F}_{\sigma_{k}}(M)$ can approximate any continuous functions on $\mathbb{B}^{d}$ if $M$ is sufficiently large. Our main theorem estimates the rate of this approximation for Hölder class.
@@
The proof of Theorem 2.1 is deferred to the next section. Our proof uses similar ideas as [3, Proposition 3], which obtained the same approximation rate for $\alpha=1$ (with an additional logarithmic factor). The conclusion is more complicated for the critical value $\alpha=(d+2k+1)/2$. We think this is due to the proof technique and conjecture that $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for all $\alpha\geq(d+2k+1)/2$, see Remark 3.4. Nevertheless, in practical applications of machine learning, the dimension $d$ is large and it is reasonable to expect that $\alpha<(d+2k+1)/2$.
@@
where we restrict the measure $\mu$ to be a discrete one supported on at most NN points. The next proposition shows that any function in $\mathcal{F}_{\sigma_{k}}(M)$ is the limit of functions in $\mathcal{F}_{\sigma_{k}}(N,M)$ as $N\to\infty$.
Proposition $2.$
For $k\in\mathbb{N}_{0}$, $\mathcal{F}_{\sigma_{k}}(1)$ is the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{\infty}(\mathbb{B}^{d})$.
@@
Let us denote the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{\infty}(\mathbb{B}^{d})$ by $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$. We first show that $\mathcal{F}_{\sigma_{k}}(1)\subseteq\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$. For any $f\in\mathcal{F}_{\sigma_{k}}(1)$ with the integral form $f(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu(v)$, we can decompose $f$ as
$f(x)$ $=\|\mu_{+}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{+}(v)}{\|\mu_{+}\|}-\|\mu_{-}\|\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)\frac{d\mu_{-}(v)}{\|\mu_{-}\|}$
@@
where $\mu_{+}$ and $\mu_{-}$ are the positive and negative parts of $\mu$. If $f_{+},f_{-}\in\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$, then $f\in\widetilde{\mathcal{F}}_{\sigma_{k}}(1)$. Hence, without loss of generality, we can assume $\mu$ is a probability measure. We are going to approximate $f$ by uniform laws of large numbers. Let $\{v_{i}\}_{i=1}^{N}$ be NN i.i.d. samples from $\mu$. By symmetrization argument (see [60, Theorem 4.10] for example), we can bound the expected approximation error by Rademacher complexity [7]:
@@
where $(\epsilon_{1},\dots,\epsilon_{N})$ is an i.i.d. sequence of Rademacher random variables. For $k\in\mathbb{N}$, the Lipschitz constant of $\sigma_{k}$ on $\sqrt{2},\sqrt{2}]$ is $k2^{(k-1)/2}$. By the contraction property of Rademacher complexity [29, Corollary 3.17],
@@
For $k=0$, the VC dimension of the function class $\{f_{x}(v)=\sigma_{0}((x^{\intercal},1)v):x\in\mathbb{B}^{d}\}$ is at most $d$ [60, Proposition 4.20]. Thus, we have the bound $\mathcal{E}_{0}(N)\lesssim\sqrt{d/N}$ by [60, Example 5.24]. Hence, $f$ is in the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$.
Next, we show that $\widetilde{\mathcal{F}}_{\sigma_{k}}(1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$ for $k\in\mathbb{N}_{0}$. Since $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)\subseteq\mathcal{F}_{\sigma_{k}}(1)$, we only need to show that $\mathcal{F}_{\sigma_{k}}(1)$ is closed in $L^{\infty}(\mathbb{B}^{d})$. Let $f_{n}(x)=\int_{\mathbb{S}^{d}}\sigma_{k}((x^{\intercal},1)v)d\mu_{n}(v)$, where $\|\mu_{n}\|\leq 1$, be a convergent sequence with limit $f\in L^{\infty}(\mathbb{B}^{d})$. It remains to show that $f\in\mathcal{F}_{\sigma_{k}}(1)$.
For $k\in\mathbb{N}$, by the compactness of $\mathbb{S}^{d}$ and Prokhorov’s theorem, there exists a weakly convergent subsequence $\mu_{n_{i}}\to\mu$. In particular, $\|\mu\|\leq 1$ and for any $x\in\mathbb{B}^{d}$,